OpenStudy (nthenic_oftime):

Use Gaussian elimination to find the complete solution to the system of equations, or state that none exists. Show your work. 4x - y + 3z = 12 x + 4y + 6z = -32 5x + 3y + 9z = 20

OpenStudy (maddielb):

Double the first equation: 8x - 2y + 6x = 24 Subtract the second equation from that: 7x - 6y = 56 Triple the first equation: 12x - 3y + 9z = 36 Subtract the third equation from that: 7x - 6y = 16 That's a contradiction. No solution.

OpenStudy (nthenic_oftime):

im sorry i dont know if i understand that... so the answer is no solution? how does doubling and tripling the equation prove that its a contradiction?

OpenStudy (nthenic_oftime):

@MaddieLB

OpenStudy (nthenic_oftime):

so by doubling the eqaution and subtracting the 2nd and the tripling and subtracting the third we should have the same result? yes?

OpenStudy (maddielb):

Yes

OpenStudy (nthenic_oftime):

is that gaussian elimination though?

OpenStudy (maddielb):

Yes, here is another look at it a different way 4 | -1 | 3 | 12 | 1 | 4 | 6 | -32 | 5 | 3 | 9 | 20 | ------------------ 1 | -0.25 | 0.75 | 3 | 0 | 4.25 | 5.25 | -35 | 0 | 4.25 | 5.25 | 5 | ------------------ 1 | -0.25 | 0.75 | 3 | 0 | 1 | 1.2353 | -8.2353 | 0 | 0 | 0 | 40 | ∅

OpenStudy (maddielb):

Alright hope that helped, I gotta go

alones (alones):

@nthenic_oftime do you understand?

OpenStudy (nthenic_oftime):

im not sure... the matricies he gave me... idk how that is relating to my equations... the first one yeah i got that but idk how the other two are the same as doubling and trippling the first one... :(

OpenStudy (nthenic_oftime):

@AloneS